# GenIAS: Generator for Instantiating Anomalies in time Series

### v1

@@ 4.5. Generation Performance
Table 3. TSAD performance of GenIAS and baseline methods using F1 score and AUPR, with the best model in bold and the second-best model underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.
@@
MTS UTS Metric Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg F1 Classic OCSVM (1999) 0.308 0.411 0.319 0.743 0.297 0.557 0.439 0.608 0.148 0.011 0.256 0.378 LOF (2000) 0.368 0.400 0.204 0.331 0.075 0.655 0.339 0.586 0.101 0.010 0.232 0.303 IForest (2008) 0.278 0.353 0.306 0.745 0.255 0.717 0.442 0.553 0.236 0.010 0.266 0.384 PU LSTM-VAE (2018) 0.407 0.437 0.298 0.724 0.208 0.625 0.450 M M M M M US-WI Donut (2018) U U U U U U U 0.489 0.126 0.009 0.208 U Omni (2019) 0.243 0.325 0.459 0.763 0.166 0.541 0.416 M M M M M THOC (2020) 0.309 0.327 0.168 0.638 0.146 0.506 0.349 0.253 0.233 0.010 0.165 0.288 MTAD-GAT(2020) 0.473 0.519 0.347 0.242 0.313 0.599 0.416 M M M M M AnomalyTran (2021) 0.345 0.407 0.304 0.738 0.155 0.524 0.412 M M M M M TranAD (2022) 0.428 0.472 0.361 0.310 0.280 0.527 0.396 0.566 0.287 0.010 0.288 0.360 TS2Vec (2022) 0.299 0.371 0.173 0.261 0.063 0.502 0.278 0.484 0.204 0.007 0.232 0.263 TimesNet (2023) 0.358 0.402 0.338 0.217 0.297 0.492 0.351 0.513 0.241 0.011 0.255 0.319 Dcdetector (2023) 0.227 0.275 0.083 0.217 0.021 0.492 0.219 0.112 0.042 0.008 0.054 0.164 US-I NCAD (2022) 0.266 0.358 0.183 0.217 0.296 0.492 0.302 0.134 0.168 0.020 0.107 0.237 CutAddPaste (2024) 0.363 0.458 0.192 0.700 0.280 0.896 0.482 0.321 0.620 0.042 0.328 0.430 COUTA (2024) 0.414 0.421 0.362 0.312 0.389 0.626 0.421 0.658 0.221 0.010 0.296 0.379 CARLA (2025) 0.523 0.529 0.511 0.721 0.293 0.908 0.581 0.749 0.308 0.039 0.365 0.509 Ours GenIAS 0.579 0.620 0.528 0.755 0.329 0.937 0.625 0.811 0.393 0.064 0.423 0.557 AUPR Classic OCSVM (1999) 0.202 0.256 0.327 0.696 0.163 0.608 0.375 0.703 0.300 0.028 0.344 0.365 LOF (2000) 0.217 0.248 0.277 0.185 0.031 0.624 0.264 0.615 0.109 0.043 0.256 0.261 IForest (2008) 0.182 0.190 0.278 0.739 0.162 0.739 0.382 0.614 0.288 0.021 0.308 0.357 PU LSTM-VAE (2018) 0.285 0.258 0.395 0.685 0.045 0.674 0.390 M M M M M US-WI Donut (2018) U U U U U U U 0.264 0.078 0.015 0.119 U Omni (2019) 0.149 0.115 0.365 0.713 0.034 0.587 0.327 M M M M M THOC (2020) 0.240 0.195 0.107 0.537 0.076 0.432 0.265 0.349 0.229 0.014 0.197 0.242 MTAD-GAT(2020) 0.335 0.339 0.401 0.095 0.200 0.613 0.331 M M M M M AnomalyTran (2021) 0.236 0.264 0.273 0.680 0.059 0.568 0.347 M M M M M TranAD (2022) 0.273 0.287 0.412 0.192 0.119 0.586 0.312 0.691 0.285 0.022 0.333 0.319 TS2Vec (2022) 0.132 0.148 0.113 0.136 0.040 0.408 0.163 0.491 0.221 0.009 0.240 0.189 TimesNet (2023) 0.283 0.208 0.385 0.082 0.189 0.508 0.276 0.671 0.237 0.032 0.313 0.288 Dcdetector (2023) 0.129 0.124 0.043 0.126 0.011 0.326 0.127 0.041 0.018 0.009 0.023 0.092 US-I NCAD (2022) 0.146 0.156 0.096 0.106 0.135 0.407 0.174 0.067 0.089 0.103 0.086 0.145 CutAddPaste (2024) 0.242 0.322 0.153 0.582 0.159 0.808 0.378 0.116 0.602 0.188 0.302 0.352 COUTA (2024) 0.247 0.248 0.400 0.159 0.264 0.685 0.334 0.555 0.282 0.049 0.295 0.321 CARLA (2025) 0.501 0.448 0.507 0.681 0.201 0.814 0.525 0.627 0.299 0.247 0.391 0.481 Ours GenIAS 0.529 0.467 0.512 0.704 0.236 0.882 0.555 0.794 0.357 0.280 0.477 0.529
4.6. Visualization
(a) Yahoo - Entity real_65
(b) SMAP - Entity E5
Figure 4. Each plot visualizes a different time series (TS) injection or generation method using t-SNE. Blue markers represent the original anomalous TS, while orange markers represent the injected or generated anomalous TS.
@@ 4.6. Visualization
To illustrate GenIAS’s enhanced anomaly generation compared to baseline methods that also leverage anomaly generation (US-I), we present t-SNE visualizations comparing their generated and real anomalies in Figure 4 on two selected datasets, Yahoo (UTS) and SMAP (MTS). Some additional visualizations are included in Figure 10 in Appendix G.2.3.
@@
Impressively, the distributions of the generated anomalies by GenIAS align with the actual anomalies distribution substantially closer compared to the baseline methods (i.e., actual anomaly regions are densely covered by the generated anomalies). This is essential to provide anomaly-discriminative information for establishing more distinctive normal and anomaly representations in unsupervised TSAD training. On the other hand, the baseline methods, to varying degrees, are unable to generate anomalies that sufficiently represent the actual anomaly distribution, which is evidenced by major blue areas (representing actual anomaly regions) that are not covered by orange areas (generated anomaly regions). For example, CARLA, the second-best method for overall ARP and TSAD performance, also suffers from these issues: on Yahoo (a), generated anomalies largely miss the anomaly regions, while on SMAP (b), generated anomalies form two major clusters leaving a gap in between, which is an actual anomaly region failing to be represented. These observations further substantiate that relying on handcrafted, predefined anomaly patterns for anomaly generation is prone to distribution misalignment when the assumed patterns deviate from actual ones, and disconnected/scattered anomaly regions when the assumed patterns differ significantly. Whereas GenIAS, through controlled and regularized perturbation on the learned distribution, generates anomalies without pattern speculation, achieving both better alignment and smoothness.
4.7. Main TSAD Performance
@@
We report the TSAD performance of GenIAS and the competing methods in terms of the best F1 score and AUPR in Table 3. Due to space limitations, please refer to the Appendix G.2.2 for results on the other three previously mentioned metrics discussed in Section 4.3. Impressively, GenIAS achieves significantly better overall TSAD performance compared to the baseline methods across all metrics, demonstrating its superior effectiveness in handling both MTS and UTS. Specifically, among methods that can operate on both types of time series, GenIAS achieves 9.4% and 29.5% performance improvements compared to the second and third best performing methods in F1, respectively. Similarly, it outperforms the second best model by 10.0% and third best model by 44.9% in terms of AUPR. Upon examining the more detailed TSAD performance with respect to factors such as method category and dataset type, we noted the following observations.
@@
Consistent effectiveness for both MTS and UTS. On each type of dataset, GenIAS yields similar levels of performance improvement as observed in overall performance. For example, compared to the second-best model for both types, GenIAS improves F1 by 7.6% and AUPR by 8.4% on MTS, and improves F1 by 15.9% and AUPR by 22.0% for UTS. This highlights its ability to generate informative and diverse anomalies that can mimic anomalous patterns within each dimension and the dependencies between dimensions. We also provide the average rank of all models in the form of critical difference diagrams (Demšar, 2006) for F1 and AUPR in both MTS and UTS (see Appendix G.2.1). GenIAS achieved the highest rank in all four comparisons and demonstrated statistically significant improvements over all baselines in MTS for both AUPR and F1.
@@
Enhanced Anomaly Generation. It is worth noting that even when compared to US-I baselines, GenIAS consistently outperforms them due to its ability to generate more realistic and diverse anomalous patterns. In particular, since GenIAS-enhanced TSAD model and CARLA only differ in their anomaly generation while sharing the same architecture and training methods, their performance gap underscores the effectiveness of GenIAS’s approach. This is attributed to its TSAD tailored variational representation learning and novel perturbation strategies, which enforce more distinctive normal representations and are not confined to predefined, handcrafted anomaly prototypes.
@@
Importance of anomaly generation in TSAD. It is notable that US-I injection methods are generally more effective than methods in other categories that do not leverage labeled anomalies. Remarkably, some of these methods, such as CutAddPaste and CARLA, can even outperform the supervised baseline LASM-VAE. This suggests that, with well-designed anomaly injection techniques, the generated pseudo anomalies can be sufficiently realistic to represent actual anomalies, including those not well captured by the labeled training data in supervised settings. Moreover, leveraging anomaly generation can generate both larger numbers and more diverse types of anomalies, which can improve TSAD representation learning. Therefore, anomaly injection should be incorporated as an essential module in TSAD models.
4.8. Ablation Study
@@
To better understand the contributions of different components and configurations in our proposed model, we conducted an ablation study. Our analysis focuses on the following aspects: (i) The impact of the patch selection, (ii) The effectiveness of the prior variance, and (iii) The effectiveness of the min-max margin.
4.8.1. Impact of the Patch Selection
First, in Figure 5, we present our study of F1 and AUPR for two patching methods—Length-driven and Deviation patching—evaluated across varying portions of patch lengths and thresholds. The x-axis represents the portion of patch length (ranging from 0.2 to 1, in intervals of 0.2), while thresholds $\tau$ are explored at 0.05, 0.1, 0.2, 0.4, and 0.6. The best results for each method, identified through this parameter exploration, are then compared in Table 4. The results in Table 4 demonstrate that Deviation patching consistently outperforms the other approaches, achieving the highest F1 and AUPR scores across all datasets. For example, it achieves a 0.044 increase in F1 and a 0.063 improvement in AUPR for SMAP compared to the baseline without patching (wo Patch). Length-driven patching, which introduces anomalies based on predefined temporal lengths, shows moderate improvements over the baseline in most scenarios but slightly underperforms on more complex datasets, such as Yahoo, where it reduces F1 and AUPR scores. The strong performance of Deviation patching, as defined in Section 3.3, underscores its ability to better align generated anomalies with true anomaly distributions, making it a robust and effective approach for anomaly generation and detection.
Table 4. Performance metrics for different patching methods.
@@ 4.8.1. Impact of the Patch Selection
Metric Dataset wo Patch Length-driven Deviation-based F1 MSL 0.535 0.547 0.579 SMAP 0.533 0.571 0.620 Yahoo 0.755 0.741 0.811 AUPR MSL 0.466 0.488 0.529 SMAP 0.425 0.429 0.467 Yahoo 0.776 0.766 0.794
Figure 5. F1 and AUPR for different patching methods.
4.8.2. Effectiveness of the Prior Variance
@@
Figure 6 analyzes the impact of prior variance on anomaly detection performance, showing consistent trends across datasets and indicating that the model is not highly sensitive to this parameter. For the best F1 scores, prior variances of 0.5 achieve optimal performance across all datasets (MSL: 0.579, SMAP: 0.62, Yahoo: 0.811), indicating a spot where the model balances bias and variance effectively. However, as the prior variance increases beyond 0.5, performance declines, suggesting that overly broad priors introduce excessive flexibility, reducing model specificity. Similarly, the AUPR results follow this trend, with 0.5 showing superior means across datasets and narrower standard deviation (std) bars, emphasizing the robustness of this setting. The increasing std with higher variances further supports that larger priors may lead to unstable and less reliable performance. Overall, a prior variance of 0.5 emerges as the most effective choice, yielding a balance between accuracy and stability for both F1 and AUPR.
Figure 6. F1 and AUPR for different prior variance ($\sigma_{\text{prior}}^{2}$)
4.8.3. Effectiveness of the Min-Max Margin
Figure 7. Heatmaps of F1 and AUPR scores across datasets illustrating the effectiveness of min-max margins.
@@ 4.8.3. Effectiveness of the Min-Max Margin
The heatmaps in Figure 7 illustrate the impact of min-max margins ($\delta_{\text{min}},\delta_{\text{max}}$) on performance metrics, such as F1 and AUPR, across MSL, SMAP, and Yahoo datasets. Smaller max margins ($\delta_{\text{max}}=0.2$) and moderate min margins ($\delta_{\text{min}}=0.1$) consistently yield the best results, with peak F1 scores of 0.579, 0.62, and 0.811 for MSL, SMAP, and Yahoo, respectively. Similarly, the AUPR values align with this trend. These results validate the design rationale in Section 3.2.3, where $\delta_{\text{min}}$ enforces distinctness, and $\delta_{\text{max}}$ maintains realism. Larger margins (e.g., $\delta_{\text{max}}=1$) reduce performance, as seen in the lower F1 scores of MSL and Yahoo. These findings emphasize the importance of balancing min-max margins.
5. Conclusion and Future Work
@@
In this work, we introduced GenIAS, a generative anomaly injection framework designed to enhance time series anomaly detection by creating realistic and diverse anomalies. By perturbing latent representations of a generative model trained on normal data, GenIAS produces anomalies that are similar to real-world anomalies. When integrated into a TSAD model, it further enhances anomaly detection performance by adding high-quality synthetic anomalies to the training set. For future work, we aim to refine adaptive perturbation techniques and an adaptive patching approach to better align with real-world domain-specific anomalies. Additionally, we plan to incorporate correlation between dimensions to guarantee the generation of intermetric anomalies, making the synthetic anomalies more representative of real-world multivariate dependencies. These enhancements will improve the realism of generated anomalies by GenIAS and its effectiveness in time series anomaly detection.
@@ Appendix B List of Symbols and Notations
Table 5. List of Symbols and Notations
@@
$p(\widetilde{\mathbf{z}})$ Shifted latent distribution for anomalies
@@ Appendix C Algorithms Overview
1Initialize encoder $q_{\phi}$, decoder $g_{\theta}$, latent space parameters $\mu$, $\sigma$, and perturbation scale $\psi$;
2 while training not converged do
      3 Sample a batch $\{\mathbf{\text{X}}_{i}\}_{i=1}^{B}\subset\mathcal{D}$ of size $B$;
      4 Encode: Obtain latent parameters $\mu_{i}$, $\sigma_{i}$ by $q_{\phi}(\mathbf{\text{X}}_{i})$;
      5 Sample latent vectors $\mathbf{\text{z}}_{i}\sim\mathcal{N}({\mu}_{i},\operatorname{diag}({\sigma}_{i}^{2}))$;
      6 Generate perturbed latent vectors $\widetilde{\mathbf{z}}_{i}$triangleright Eq. (1)
      7 Decode reconstructed samples: $\hat{\mathbf{X}}_{i}=g_{\theta}(\mathbf{z}_{i})$;
      8 Decode anomalous samples: $\hat{\mathbf{X}}_{i}\ =g_{\theta}(\widetilde{\mathbf{z}}_{i})$;
      9 Compute the reconstruction loss $\mathcal{L}_{\text{recon}}$; $\triangleright$ Eq. (2)
      10 Compute the perturbation loss $\mathcal{L}_{\text{perturb}}$; $\triangleright$ Eq. (3)
      11 Compute the zero-perturb loss $\mathcal{L}_{\text{zero-perturb}}$; $\triangleright$ Eq. (4)
      12 Compute the enhanced KL loss $\mathcal{L}_{\text{en-KL}}$; $\triangleright$ Eq. (5)
      13 $\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{en-KL}}$;
      14 Update the model parameters $\phi$, $\theta$ and $\psi$  by minimizing $\mathcal{L}_{\text{total}}$;
15return $\phi$, $\theta$, $\psi$;
Algorithm 1 End-to-End Training of GenIAS
@@ 14 Update the model parameters $\phi$, $\theta$ and $\psi$  by minimizing $\mathcal{L}_{\text{total}}$;
1 Load encoder $q_{\phi}$, decoder $g_{\theta}$, and perturbation scales $\psi$.
2Encode: Obtain latent parameters $\mu$, $\sigma$ by $q_{\phi}(\mathbf{\text{X}})$;
3Generate perturbed latent vectors $\widetilde{\mathbf{z}}$triangleright Eq. (1)
4Decode $\widetilde{\mathbf{z}}$ to generate preliminary anomalies: $\tilde{\mathbf{X}}=g_{\theta}(\tilde{\mathbf{z}})$;
@@
5 foreach dimension $d$ in $\mathbf{X}$ do
      6 $\text{deviation}_{d}=\left|\mathbf{X}_{d}-\tilde{\mathbf{X}}_{d}\right|^{2}$;
      7 $\text{amplitude}_{d}=\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d})$;
      8 Generate patched output for dimension $d$:
@@
9return $\tilde{\mathbf{X}}_{\text{patched}}$ as the patched anomalous TS windows.
Algorithm 2 Patching with GenIAS

### v2

@@ 4.6 Visualization
Table 3: TSAD performance of GenIAS and baseline methods using F1 score, with the best model in bold and the second-best model underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.
@@
MTS UTS Cat. Methods MSL SMAP SMD SWAT GECCO SWAN Avg-M Yahoo KPI UCR Avg-U Avg Classic OCSVM (1999) 0.308 0.411 0.319 0.743 0.297 0.557 0.439 0.608 0.148 0.011 0.256 0.378 LOF (2000) 0.368 0.400 0.204 0.331 0.075 0.655 0.339 0.586 0.101 0.010 0.232 0.303 IForest (2008) 0.278 0.353 0.306 0.745 0.255 0.717 0.442 0.553 0.236 0.010 0.266 0.384 PU LSTM-VAE (2018) 0.407 0.437 0.298 0.724 0.208 0.625 0.450 M M M M M US-WI Donut (2018) U U U U U U U 0.489 0.126 0.009 0.208 U Omni (2019) 0.243 0.325 0.459 0.763 0.166 0.541 0.416 M M M M M THOC (2020) 0.309 0.327 0.168 0.638 0.146 0.506 0.349 0.253 0.233 0.010 0.165 0.288 MTAD-GAT (2020) 0.473 0.519 0.347 0.242 0.313 0.599 0.416 M M M M M AnomalyTran (2021) 0.345 0.407 0.304 0.738 0.155 0.524 0.412 M M M M M TranAD (2022) 0.428 0.472 0.361 0.310 0.280 0.527 0.396 0.566 0.287 0.010 0.288 0.360 TS2Vec (2022) 0.299 0.371 0.173 0.261 0.063 0.502 0.278 0.484 0.204 0.007 0.232 0.263 TimesNet (2023) 0.358 0.402 0.338 0.217 0.297 0.492 0.351 0.513 0.241 0.011 0.255 0.319 Dcdetector (2023) 0.227 0.275 0.083 0.217 0.021 0.492 0.219 0.112 0.042 0.008 0.054 0.164 US-I NCAD (2022) 0.266 0.358 0.183 0.217 0.296 0.492 0.302 0.134 0.168 0.020 0.107 0.237 CutAddPaste (2024) 0.363 0.458 0.192 0.700 0.280 0.896 0.482 0.321 0.620 0.042 0.328 0.430 COUTA (2024) 0.414 0.421 0.362 0.312 0.389 0.626 0.421 0.658 0.221 0.010 0.296 0.379 CARLA (2025) 0.523 0.529 0.511 0.721 0.293 0.908 0.581 0.749 0.308 0.039 0.365 0.509 Ours GenIAS+CARLA 0.579 0.620 0.528 0.755 0.329 0.937 0.625 0.811 0.393 0.064 0.423 0.557
@@
To illustrate GenIAS’s enhanced anomaly generation compared to other anomaly injection-based baselines (US-I), we present t-SNE visualizations in Figure 4, comparing generated and real anomalies on Yahoo. Additional visualizations are provided in Figure 10 in Appendix G.2.3.
@@
Impressively, GenIAS-generated anomalies align substantially more closely with the actual anomaly distribution compared to baseline methods, effectively covering true anomaly regions. This is essential to provide anomaly-discriminative information for establishing more distinctive normal and anomaly representations in unsupervised TSAD training. In contrast, baseline methods struggle to generate anomalies that adequately capture the actual anomaly distribution, as shown by large blue regions (actual anomalies) left uncovered by orange regions (generated anomalies). For example, CARLA, the second-best method in overall ARP and TSAD performance, struggles on Yahoo, where its generated anomalies largely miss true anomaly regions. This highlights a key limitation of relying on handcrafted, predefined anomaly patterns: they risk distribution misalignment when assumptions deviate from reality, often resulting in disconnected or scattered anomaly regions. In contrast, GenIAS leverages controlled, regularized perturbation on the learned distribution to generate anomalies without pattern speculation, achieving better alignment and smoothness.
4.7 Main TSAD Performance
@@
We report the TSAD performance of GenIAS and the competing methods in terms of the best F1 score in Table 3. Due to space constraints, results for the other four metrics discussed in Section 4.3 are provided in Appendix G.2.2. GenIAS achieves significantly better overall TSAD performance compared to the baselines across metrics, demonstrating its superior effectiveness on both MTS and UTS. Specifically, among methods applicable to both types of time series, GenIAS achieves F1 improvements of 9.4% and 29.5% over the second- and third-best, respectively. A more detailed analysis by method category and dataset type reveals the following key observations.
@@
Consistent effectiveness for both MTS and UTS. On each type of dataset, GenIAS yields similar levels of performance improvement as observed in overall performance. For example, compared to the second-best model for both types, GenIAS improves F1 by 7.6% on MTS and improves F1 by 15.9%. This highlights its ability to generate informative and diverse anomalies that can mimic anomalous patterns within each dimension and the dependencies between dimensions. We also provide the average rank of all models in the form of critical difference diagrams (Demšar, 2006) for F1 and AUPR in both MTS and UTS (see Appendix G.2.1). GenIAS achieved the highest rank in all four comparisons and demonstrated statistically significant improvements over all baselines in MTS for the F1 score.
@@
Enhanced Anomaly Generation. GenIAS consistently outperforms US-I baselines by generating more realistic and diverse anomalous patterns. In particular, its performance gain over CARLA while only differing in their anomaly generation highlights the effectiveness of GenIAS’s tailored variational learning and novel perturbation strategies, which enforce more distinctive normal representations without relying on handcrafted anomaly types.
@@
Importance of anomaly generation in TSAD. US-I injection methods generally outperform other unsupervised approaches and, in some cases, like CutAddPaste and CARLA, can even outperform the supervised baseline LSTM-VAE. This suggests that well-designed anomaly generation can produce realistic pseudo-anomalies that capture true anomalies, including those missed by labeled data. Moreover, it enables greater anomaly diversity, improving representation learning. Therefore, anomaly injection should be considered as an essential module in TSAD models.
4.8 Ablation Study
@@
To better understand the contributions of different components and configurations in our proposed model, we conducted an ablation study.
Table 4: Performance comparison of patching methods
@@
Dataset wo Patch Length-driven Deviation-based F1 AUPR F1 AUPR F1 AUPR MSL 0.535 0.466 0.547 0.488 0.579 0.529 SMAP 0.533 0.425 0.571 0.429 0.620 0.467 Yahoo 0.755 0.776 0.741 0.766 0.811 0.794
Figure 5: F1 and AUPR for different patching methods.
Impact of the Patch Selection. Figure 5 presents our analysis of F1 and AUPR for two patching methods—Length-driven and Deviation patching—evaluated across varying patch lengths and thresholds. The x-axis indicates the varying portion of patch length and the threshold $\tau$ values for these approaches. The best results from this parameter study are compared in Table 4, demonstrating that Deviation patching consistently outperforms other approaches, achieving the highest F1 and AUPR scores across all datasets. For example, on SMAP, it improves F1 by 0.044 and AUPR by 0.063 compared to the baseline without patching (wo Patch). Length-driven patching, which injects anomalies based on predefined temporal lengths, shows moderate improvements over the baseline in most scenarios but slightly underperforms on Yahoo dataset, where it reduces F1 and AUPR scores. In contrast, the strong performance of Deviation patching (Section 3.3) highlights its effectiveness in aligning generated anomalies with true anomaly distributions, making it a more robust and effective approach for anomaly generation and detection.
Figure 6: F1 and AUPR for different prior variance ($\sigma_{\text{prior}}^{2}$)
@@
Effectiveness of the Prior Variance. Figure 6 analyzes the impact of prior variance on anomaly detection performance, showing consistent trends across datasets and low sensitivity to this parameter. A prior variance of 0.5 yields the best F1 scores across all datasets (MSL: 0.579, SMAP: 0.62, Yahoo: 0.811), suggesting an effective balance between bias and variance. However, as the prior variance increases beyond 0.5, performance declines, indicating that overly broad priors reduce model specificity. Similarly, the AUPR results follow this trend, with 0.5 showing superior means and the smallest standard deviation (std), emphasizing the robustness of this setting. The growing std at higher variances further supports that large priors can lead to unstable and less reliable performance. Overall, a prior variance of 0.5 emerges as the most effective choice, yielding a balance between accuracy and stability for both F1 and AUPR.
Figure 7: Effectiveness of min-max margins.
@@
Effectiveness of the Min-Max Margin. The heatmaps in Figure 7 illustrate the impact of min-max margins ($\delta_{\text{min}},\delta_{\text{max}}$) on F1 across MSL, SMAP, and Yahoo datasets. Smaller max margins ($\delta_{\text{max}}=0.2$) and moderate min margins ($\delta_{\text{min}}=0.1$) consistently yield the best results, with peak F1 scores of 0.579, 0.62, and 0.811 for MSL, SMAP, and Yahoo, respectively. These results validate the design rationale in Section 3.2, where $\delta_{\text{min}}$ enforces distinctness and $\delta_{\text{max}}$ maintains realism. Larger margins (e.g., $\delta_{\text{max}}=1$) reduce performance, as seen in lower F1 scores of MSL and Yahoo, emphasizing the importance of balancing margins.
5 Conclusion and Future Work
@@
We introduced GenIAS, a generative anomaly injection framework designed to enhance time series anomaly detection by creating realistic and diverse anomalies. By perturbing latent representations of a generative model trained on normal data, GenIAS produces anomalies that are similar to real-world anomalies. Integrated into a TSAD model, it boosts anomaly detection performance by enriching training data with high-quality synthetic anomalies.
For future work, we will explore adaptive perturbation and patching strategies, along with modelling inter-dimensional correlations, to further improve synthetic anomaly generation for TSAD.
Impact Statement
GenIAS improves TSAD by enabling realistic and controllable synthetic anomaly generation using variational neural models. This alleviates the difficulty of training robust algorithms in application domains with few labeled anomalies, particularly for complex multivariate time series with subtle anomalous patterns.
@@ Appendix B List of Symbols and Notations
Table 5: List of Symbols and Notations
@@
[…]
@@ Appendix C Algorithms Overview
Initialize encoder $q_{\phi}$, decoder $g_{\theta}$, latent space parameters $\mu$, $\sigma$, and perturbation scale $\psi$
while training not converged do
 14131211109876543 Sample a batch $\{\mathbf{\text{X}}_{i}\}_{i=1}^{B}\subset\mathcal{D}$ of size $B$  Encode: Obtain latent parameters $\mu_{i}$, $\sigma_{i}$ by $q_{\phi}(\mathbf{\text{X}}_{i})$  Sample latent vectors $\mathbf{\text{z}}_{i}\sim\mathcal{N}({\mu}_{i},\operatorname{diag}({\sigma}_{i}^{2}))$  Generate perturbed latent vectors $\widetilde{\mathbf{z}}_{i}$ $\triangleright$ Eq. (1)
Decode reconstructed samples: $\hat{\mathbf{X}}_{i}=g_{\theta}(\mathbf{z}_{i})$  Decode anomalous samples: $\hat{\mathbf{X}}_{i}\ =g_{\theta}(\widetilde{\mathbf{z}}_{i})$  Compute the reconstruction loss $\mathcal{L}_{\text{recon}}$; $\triangleright$ Eq. (5)
Compute the perturbation loss $\mathcal{L}_{\text{perturb}}$; $\triangleright$ Eq. (3)
Compute the zero-perturb loss $\mathcal{L}_{\text{zero-perturb}}$; $\triangleright$ Eq. (4)
Compute the compact KL loss $\mathcal{L}_{\text{comp-KL}}$; $\triangleright$ Eq. (2)
$\mathcal{L}_{\text{total}}=\alpha\cdot\mathcal{L}_{\text{recon}}+\beta\cdot\mathcal{L}_{\text{perturb}}+\gamma\cdot\mathcal{L}_{\text{zero-perturb}}+\zeta\cdot\mathcal{L}_{\text{comp-KL}}$;
Update the model parameters $\phi$, $\theta$ and $\psi$ by minimizing $\mathcal{L}_{\text{total}}$
15return $\phi$, $\theta$, $\psi$
Algorithm 2 Patching with GenIAS
@@
Load encoder $q_{\phi}$, decoder $g_{\theta}$, and perturbation scales $\psi$.
Encode: Obtain latent parameters $\mu$, $\sigma$ by $q_{\phi}(\mathbf{\text{X}})$
Generate perturbed latent vectors $\widetilde{\mathbf{z}} \triangleright$ Eq. (1)
[…]
@@
foreach dimension $d$ in $\mathbf{X}$ do
 232221 $\text{deviation}_{d}=\left|\mathbf{X}_{d}-\tilde{\mathbf{X}}_{d}\right|^{2}$$\text{amplitude}_{d}=\max(\mathbf{X}_{d})-\min(\mathbf{X}_{d})$  Generate patched output for dimension $d$:
@@
24return $\tilde{\mathbf{X}}_{\text{patched}}$ as the patched anomalous TS windows.